OpenStudy (anonymous):

Find the derivative of:

OpenStudy (anonymous):

\[7\div-3\cos 3x\]

OpenStudy (jhannybean):

\[\frac{7}{-3\cos{3x}}\]Remember, a constant can always be factored out, or you could use the quotient rule.

OpenStudy (anonymous):

When I worked it out, i ended up with an extra number

rvc (rvc):

can we make this sec3x and then diff @rational

OpenStudy (anonymous):

\[7\sin3x \div-3\cos^23x\] this is what i got

OpenStudy (anonymous):

we havent learnt sec

OpenStudy (jhannybean):

So... \[\frac{\frac{d}{dx}(7)(-3\cos(3x))-\frac{d}{dx}(-3\cos(3x))(7)}{(-3\cos(3x))^2}\]

OpenStudy (jhannybean):

If you're just getting used to taking derivatives.

OpenStudy (kainui):

A quick derivation of the derivative of sec x if you want me to explain any part of this more feel free to ask :) \[\large \frac{d}{dx}\sec(x) = \frac{d}{dx} (\cos x )^{-1}=-1(\cos x)^{-2}(-\sin x)=\tan x \sec x\]

OpenStudy (anonymous):

I think I figured out what i did wrong. Thanks anyway for the help!!

OpenStudy (jhannybean):

Which method did you use?

OpenStudy (jhannybean):

You could also use the product rule which @Kainui hinted at. \[7(-3\cos(3x))^{-1}\]where \(f(x) =7\) and \(g(x) = -3\cos(3x)\)

rvc (rvc):

anyways u got ur mistake and now its cleared All the best @adriana_98

OpenStudy (anonymous):

Thanks. Exam is tomorrow :(

OpenStudy (jhannybean):

Good luck. May Newton be with you.

rvc (rvc):

lol

OpenStudy (anonymous):

Thanks!! @Jhannybean

OpenStudy (jhannybean):

No problem :)